D. A. H. Buckley M. A. Bershady

Sizin üçün oyun:

Google Play'də əldə edin


Yüklə 77.1 Kb.
tarix26.07.2018
ölçüsü77.1 Kb.


Minutes of the 13th SALT Science Working Group meeting
8 & 9 May 2005
SALT Boardroom, SAAO, Cape Town


D.A.H. Buckley

M. A. Bershady



20 May 2005


1.Participants

Members:


Matt Bershady (Proxy: Wisconsin)

Stuart Barnes (Proxy: U. Canterbury)

Gordon Bromage (Proxy: UK Consortium)

David Buckley (Project Scientist, Chair)

Brian Chaboyer (Dartmouth College)

Richard Griffiths (CMU)

Janusz Kaluzny (Poland)

Wolfram Kollatschny (Göttingen)

Ken Nordsieck (Wisconsin-Madison)

Darragh O’Donoghue (South Africa)


Larry Ramsey (HET)

Ted Williams (Rutgers)


Ex-officio attendees:


Phil Charles (SAAO Director)

Peter Cottrell (Prox: U. Canterbury)

Roy Emmerich (SALT)

Michael Feast (UCT)

Robert Fesen (Dartmouth College)

Kobus Meiring (SALT)

Marek Sarna (Poland)

Peter Sarre (Nottingham)

Arfon Smith (Nottingham)

Gerhard Swart (SALT)

Jian Swiegers (SALT)

Hannah Wouters (UCLAN)





2.Welcome and Minutes of the previous SSWG meeting


The participants were welcomed by the Project Scientist.
The minutes of the 12th SSWG meeting (May 2004) were tabled. The only corrections regarded the Rutgers report, where the number of clusters in the survey should be changed from 100 to 1000, and that Larry Ramsey was indeed present for the meeting (left of participants list). The corrected minutes were unanimously carried (moved Larry, seconded Ted).

3.Agenda


Monday 8th May

  1. SALT Schedule, Completion, Commissioning & Testing (Kobus) (08:30 – 09:00)

  2. SALT System Engineering, TCS and optimization (Gerhard) (09:00 – 09:45)

  3. SAMS System Update (Jian) (09:45 – 11:00)

  4. Coffee/Tea (11:00 – 11:15)

  5. Update on Planning Tools, Scheduling, etc. (Roy/Encarni) (11:20 – 12:00)

  6. PV phase status (David / Phil) (12:00 – 12:30)

  7. Report on SALT astronomy operations (David / Phil) (12:30 – 13:00)

  8. Lunch (13:00 – 14:00)

  9. SALTICAM & CCD Detector Status (Darragh) (14:00 – 14:20)

  10. SALTICAM ACSI report (Darragh) (14:20 – 14:40)

  11. PFIS Status Report (Ken) (14:40 – 15:30)

12. Coffee (15:30 – 16:00)

13. PFIS Near IR Beam concept telecon discussion (Andy Sheinis) (16:00 – 17:00)
Tuesday 9th May

14. Report on HRS CDR (David) (08:30 – 08:50)

15. HRS Status report (Peter/Stuart Barnes) (08:50 – 09:30)

16. SALT Future Development & 2nd Gen. Instruments (09:30 – 10:30

(Darragh + everyone)



17. Coffee (10:30 – 10:45)

19. SALT Partner reports (everyone) (10:45 – 11:15)

18. First Science, First Light and Publicity Discussions (everyone) (11:15 – 12:30)

19. Other business (12:30 – 13:00)

4.SALT Schedule, Completion, Commissioning & Testing: JGM


Kobus explained the problems associated with observing on Sat night during the visit to SALT:
* Dome not conditioned -5 deg C delta opening.

* Cold igloo left open - laser interferometer at 9 deg instead

of 20 deg C (tolerance +/- 0.5 deg).

* New mirrors not properly aligned

* Poor coordination - no operations personnel present over the weekend

* Mirror alignment process slow

* Tracker autofocus not functioning properly because of interferometer

at wrong temp

* Bad stacking, poor focusing
SOLUTION: Hand over SALT to SAAO for day-to-day operations, focus SALT team effort on the remaining work.
Work break-down structure: items left to do include the following:
1. Tracker

Fit shock absorber and extra skew sensor - Leon, June

RRS "list C" items - Leon Jun-Aug
2. Payload

Fit the ADC, moving baffle, guiders - Leon, Jun

Complete and fit calibration system - Willem, Jun


  • This unit is still partially under design; some components manufactured.

ATP's on above - Leon, Jun

3. SALTICAM autoguider

James effort for period May-Jul.

Design review is still to be done; bearings and motors ordered

4. Primary Mirror

Complete SAMS installation - Jian, Jun-Aug

Improve SAMS performance - Jian, Jun-Aug

Install mirror cleaning - Willem, Jun
5. TCS:

PFIS integration, PIPT, Science Data Base - Gerhard, Jun-Dec


6. System level:

Testing, Mainly IQ and environment - Gerhard, Jun-Aug, thereafter SALT ops.

Final VCRM rev (<20%) - Kobus, Aug
7. VCRM Status:

Image-quality and environmental – ongoing

Pointing & tracking – being done
The handover strategy from the Project Team to SALT Operations was presented by Phil and discussed. The following transfer of responsibilities would take place:

From To Consultant Basis

Kobus Meiring Phil Charles Meiring 20-80% for 3 months

Mike Lomberg Faranah Osman Lomberg 25% for 7 months

Gerhard Swart Herman Kriel Swart 100% for 1 month tapering

down to 20% by Dec

Jian Swiegers Hitesh Gajjar Swiegers 20-80% for 4 months

Leon Nel replacement Nel 20-50% for 3 months;

TBD 20% for following 9 months

Willem Esterhuyse Mark Gordon Esterhuyse 20-50% for 3 months

Arek Swat O'Donoghue Swat no physical presence after Jun

Clifford Gumede Johan Scholtz Swart Ad hoc
Summary:

Facility structure and dome are performing well.

Tracker performing well, small iusses outstanding reliability is the issue; performance otherwise very good. Payload ready to be installed. Primary mirror segments all installed.
SAMS installed (37out 91 segments), testing ongoing.
Basic TCS ok, enhanced TCS won’t be ready until year end.
VCRM 80-90% complete by end May.
Operations team are trained and ready, but SALT site systems engineer and Vusi's (Tracker Technician) replacement's appointment critical. Some outstanding contractor training.
SALT team will hand over to SAAO by end of May.
SALT Team will stay part-time to complete specific work and ensure smooth handover.
Discussion:

3 SALT operators have been appointed and are in training, 4th appointed to start in July.

SALT Astronomers: 1 in place; 2 others appointed.

SALT: transition to operations

This cannot wait until SALT completed

Ops team must start to obtain experience

Helps test true status of SALT
Problems:

Recognised need for SALT Systems Engineer (Gerhard’s replacement) since JOC was occupied with general S/L ops and other engineering tasks. The resignation of Vusi has hightened the need for additional staff and staff amenities need to set a target date for handover of ops responsibility to end May.


Early Apr: started seeking new Systems Engineer and decided to focus on completing key components for end May

-> Gerhard put in charge of all SALT engineering activities (both Project and Ops)


Phil mentioned that the new site Systems Engineer, Herman Kriel, has accepted an offer and will begin in about 1 month. He will be resident in Sutherland. A revised organogram for SALT operations was also shown.
Ops Structure:

PAC: Director

JOC: Sutherland site manager

Gerhard Swart: SALT Eng Ops

DB: SALT astronomy ops

Dave Kilkenny: SAAO small telescopes


MAJOR ISSUE: JOC’s time MUST be freed up to devote full time to complete SALTICAM autoguider.
MAB, RF and TW all voiced concern that JOC time must be freed up; reporting structure within the new transition structure must be carefully defined.
KN asked for a list of ops personnel.

5.SALT System Engineering: GS (Gerhard Swart)


Tracker:

Interruptions per track now below 1 in 10 since Mar'05. The PMAC motion control card also now seems to be more reliable. Two primary fault causes now: premature envelop limit; while slewing, an intermittent error reported from tracker which TCS catches and stops slew. Both are software-induced within LabVIEW, and can be fixed fairly straightforwardly. Time lost ought to be minimal?


Pointing accuracy: has decreased well below 100 arcsec, with a nice steady decrease to present value of ~17 arcsec (spec is 15 arcsec).
Image quality: again, nice smooth trend from 5 arcsec EE50 on Sep'03 to 1.3 arcsec EE50.
IQ test procedure:

1. Quantify performance of DIMMS

2. Quantify measurement accuracy

3. Determine sample size and pass/fail criteria

4. Test in available configurations:


  • bare CCD at PFIS focal station

  • Hartmann camera at aux port

  • SALTICAM ACSI

5. Take into account items influencing results

  • seeing

  • ADC in/out

  • mirror alignment (edge sensors active / in-active)

  • accuracy of measurement device

6. calculate weigted IQ as function of field position


On-sky images look promising (more testing required)
Individual error contributors are within spec:

+ mirror segment/mount excellent

+ mirror alignment good

+ tracker positioning within spec

+ tracker tracking to within spec

+ structure movement under wind force within spec

+ SAC within spec

+ guidance and focus looks okay (more testing required)


Retention of IQ depends on SAMS performance

+ SAMS improves temp. range but long-term work required

+ trade-off alignment retention vs time-to-align
Sample data with bare CCD camera (“Clemens camera”):

58 arcsec FOV

76 segments

2h after alignment to 0.06 arcsec specification

No SAMS, no ADC

0.2 C temp change

After 40 min tracking

Interferometer off

seeing not measured at the time

Results of on-sky IQ: 1.07 arcsec FWHM, 1.35 arcsec EE50, 2.58 arcsec EE80

10 sec exposure

coma visible in image - tip-tilt error possible, or perhaps some segments mis-aligned

no filter - Apogee camera, blue sensitive
Data products:

Science Data Base (SDB) record of actual observation parameters, including path to FITS data. FITS data (include observation-specific calibration) and general calibration data (ASCII), if requested:

Actual tracker trajectory vs time

Actual pupil size vs time

Environment conditions vs time

Operator log for entire night

Astronomer log for specific observation

Acquisition image to confirm pointing


TCS: Typical observation process

Observations planned using PIPT:

Data is sent to science database via XML, parsed at database

OPT (in SAMMI) reads SDB data and SA schedules observation

SOMMI reads data and SO stars observation process

TCSS (server) commands telescope subsystems to required position. TCSS reads data and sends configuration commands to instrument, based on details in SDB.

SO adjusts pointing and stars guidance.

SA checks instrument configuration and stars exposure.

FITS data saved on detector computer.

Also immediately sent o QCPC (Quality Control PC) via 1Gb/s LAN

Data is copied from there to DPC computer (RAID 5 disks)

Data backed-up daily to DVD and Cape Town


ACTION ITEM-1 (GS): update and distribute error budget table.

6.SAMS System Update (Jian Swiegers)


The installation schedule was presented. Delivery of components is currently on schedule and SAAO staff are trained to independently bond sensors. ~1 segment/day can be bonded.
sub-item % completion

plates 41%

sensors 36

cables 75

small cables 51

racks 50


capacitive modules 38

reference sensors 100%


SAMS wiring is involved and still a lot more to do. SAMS sensors are scanned (with digital imaging scanner, as a final check for imperfections and to keep a record.
Robustness: FOGALE is undertaking high-humidity testing.

De-lamination problems: nickel-gold layer now not used and parylene (dielectric) is applied to good quality Cu sensor plates. This appears to be working well.


In-situ testing: John Menzies and Hitesh Gajjar have worked 2 weeks with Alain Courterville from FOGALE. They detected components not functioning correctly, and replaced them, as well as improving the connector design (for small cables). A slow drift is sometimes seen due to some cables, capacitive modules, and racks (these to be replaced).
Calibrate at low relative humidity (RH) for dT (dGap): two independent methods:

  • minimize rms matrix error

  • test TTP maintenance using CCAS

SAMS racks are kept at dT = 0.2 to 0.4 C.


Goal of 50 nm noise per sensor. If so, ability to align to spec should be possible. Approach to constraining the matrix inversion is still under development. Fixing single mirror in all d.o.f. not as good as fixing triad in one d.o.f. Tracker model, however, based on assumption of fixing single mirror. Effect should be tracker error; will fold into budget and see if substantial. Because all mirror actuators are encoded, this should be correctable in software, although mirror

changes will require re-calibration.


A nice demonstration of the matrix-inversion alignment tool was shown by Jian.
A set of diagnostic tools to evaluate hardware problems now exists. These help to isolate bad sensors and cables via Figure of Merit (FoM). The FoM tracks with humidity, more than expected below 65% RH. It appears to be related to ability to ground front surface of mirror – it needs to be better than 10 ohms. Mirror gaps tracks with temp (as expected). Nice examples of real data were shown, taken in different delta temp, RH, etc. Can detect person climbing on truss, but returns to previous so no hysteresis.
According to simulator, we might expect ~75 nanometers noise per segment.
Brian Chaboyer asked about who at SAAO is learning about how to use these tools and interpret them. John Menzies and Hitesh Gajjar are primarily involved and are now well experienced.
Summary of performance:

After removing problems, performance went from 1500 nm/C to 200 nm/C and FoM from 200 nm/C to 70 nm/C after optimizing. Average over 100 readings when referenced; RMS FoM starts at 20 nm (previously 50nm). Retain GroC to with 0.2 deg. Best alignment: 0.1 arcsec for dT = 1.4 deg C.


Next steps:

  • fix grounding problems

  • add next ring

  • increase involvement of John Menzies and Lucian Botha (SALT Operator)

  • study temperature and RH history on site to evaluate effect of not meeting 20deg C and 92% RH spec

  • create real-time logger/analyzer using current logger & tools as basis. To be used in final SAMS systems, and will used various criteria to detect problems, supply management/maintenance information, assist with calibration (perhaps during closed loop control?). Requirements document will be compiled (HG, JM, LB)

  • continue with calibration runs

  • find calibration methods that limits night-time requirement – this is part of the delivery process from the Project Team.

  • reference sensors: solve grounding. Use to calibration RH.

  • evaluate different constraints

  • mount transfer function improved

  • loop gain currently 1: test effect damping or filtering.

  • integration of in-plane deformations to compute more accurate A-matrix coeff.

  • use true gap distance instead of average.

  • calculate in-plane mirror deformation based on gap measurements

  • in situ measurement of A-matrix by moving actuators and measuring sensor response.

  • better use of FoM, and possibility of using daytime measurements.

PMS IQ Error Budget: total 0.439” (spec); 0.436” (actual). There is ONE sub-component that is out of spec: alignment maintenance. If this could be brought to spec, global PMS performance would be significantly better than spec. The key thing for mirror alignment is low wind and < 5 deg temp variance.


Hardware reliability:

Grounding problems: 3M copper tape from ion-figuring point on mirror front-face to back. First batch ok, but new batch has few Kohm resistance. Small cables (white): strict spec on weight on segment. When installing cable, they sometimestwist in connection action and sometimes break or are damaged.


Over past two weeks, hardware is fairly robust. Once bad components identified, not case of subsequent failure. Not a good sense yet of how much hardware failure in grounding and small cabling will occur, but tractable problem.

7.Update of Planning Tools - Roy Emmrich


A brief verbal update (no presentation) was given. Roy’s new job as SAAO IT manager has meant not so much time as he’d hoped to work on the PIPT. This is a resource allocation problem.
KHN noted that is was a concern for PFIS PI tools since Jun was time for PFIS to start their development -- a time when there would be little support available from SAAO.
PC noted this was an issue of not having enough staff on board at this time.

8.Telescope Efficiencies (DB)


David reviewed, for information, the general specified SALT efficiencies, including the acquisition time spec.
Less than 9 min total acquisition time (from start of slew command to final closed loop guidance and positioning of object on slit/fibre).
Comparison of HET/SALT HET SALT

CCD exposure and readout 38% 66%

slew, acquire and guide 25% 20%

Instr. Calib. 3% 5%

PMS alignment 25% 5%

Science percentage 26% 50% (these compare to Keck, VLT, Gemini: 43-62%)



9.PV Phase Discussion (DB)


Current schedule:

15 May 05 Phase-I for SALTICAM (with no auto-guider – rely on open-loop performance)

15 Jun 05 Phase-I for PFIS

15 Jun 05 PI informed on outcome of SALTICAM PV

01 Jul 05 Beginning PV phase for SALTICAM

01 Aug 05 PIs informed on outcome of PFIS PV

01 Sep 05 Beginning of SALTICAM PV phase
Discussion:

DOD: proposals will be assessed on "can we do this now" not on scientific merit. Technical merit will be much more crucial. Gerhard Swart: he will get open-loop tracking numbers to us. Better in south than north (Dec dependent). TW: Drift-scanning will be available at some point in SALTICAM PV-phase. DB: Let's push back Phase-I until 30 May 05 for SALTICAM. KN: Doesn't want PV deadline to slip for PFIS. PC and DB say "let's be real" 30 Jun 05.


Summary of revised plan:

30 May 05 SALTICAM PV-phase deadline

30 Jun 05 PFIS PV-phase deadline
DOD SALTICAM: 8 slot wheel UBVRI+clear+ ND.

Two UV narrow band filters:

Centre wl FWHM LyB MgII OII

340 nm 40 nm 1.796 0.215

380 nm 40 nm 2.125 0.020 (0.0142-0.0250)
The response function is fairly square; TW has the manufacturer's transmission plot and will pass

this onto DoD, who will scan and put it on the web site.


On the fly free-wheeling discussion results in this schedule as noted by DB:
01 Jul 05 - 10 Nov 05 SALTICAM PV

01 Sep 05 - 31 Dec 05 PFIS PV

15 Sep 05 Phase-1 for First-half semester

15 Nov 05 Phase-2 for First-half semester

01 Jan 06 - 31 Mar 06 First half-semester normal operations

01 Apr 06 - 31 Sep 06 First full semester normal operations


Issue of how much time is really needed between phase-1 and start of normal ops semester. LWR thinks its too long.

10.Report on Astronomy Operations (DB)


Three SALT Astronomers (SAs) will be in place as of 01 Jul, but are available now "as needed".

Astronomy Ops. team: total number of SA people is 6, with total FTE of 4.2. SA: Encarni available now and one other, Petri Vaisanen, is also available.


3rd person offered SA position: Martin Still from Goddard (SWIFT and XMM, X-ray and ground-based optical). Clearly a top person.
The remaining 3 SA candidates have not yet been interviewed, but the calibre of applicants is very high. Obviously looking for people with postdoctoral, observational, software and instrument experience. When hired? As fast as SAAO can.

11.SALTICAM ACSI - Status Report (DOD)


Optical assembly: $100K, JOC and Bruce Bigelow (BB) - optomechanics

BB assembled into cells and cells assembled and aligned by DoD into the barrel.

Optical testing: original estimate: 1 day, DoD updated to 1 week; and actual was 1 month (c’est la vie).
Testing has been done using a wavefront camera (from Oko technology, about $4K).

Optical testing summary: rms wavefront of 0.25-0.4 wave

Results in EE50 of 0.17-0.225 (spec 0.3) and EE80 of 0.35 (spec is 0.5).

The testing is only on axis and at one wavelength.


Still to do:

On-telescope tests of real on-telescope IQ vs field, wavelength, temp., rho-stage position, and focus actuation.


Needed better engineering on test-mirror to establish tip-tilt (focus ok).
Residual peal-valley variation of barrel w.r.t to instrument frame of 120 microns which gives 0.5 arcsec defocus at extremes of motion.
Schedule:

About a year late (along with telescope) though VI was installed ~20 months ago.

Why the delay :


  • inexperience of PI

  • near disappearance of key staff (JOC and Faranah Osman)

  • overall project schedule was very aggressive


Money:

  • SAAO has observed all of the staff time over-run, incl. L. Balona’s (budgeted).

  • SAAO control came in below PDR price.

  • The delay has not costed project money.


Next:

Ghost investigations. Telescope CCD detector monitoring.

At time of auto-guider installation:


  • barrell vs framework flexure fix

  • improve light-tightness (no duct tape)

  • 3rd version of vacuum valve

  • science functionality and PI support

Confident that it is ready as an acquisition camera, but not yet as a science imager. VI is now decommissioned and.ACSI was “shipped” today! The money is OK and the key thing now is to improve quality.


Detector:

Luis Balona went over various detector issues including the CCD characterization, bias measurements, pre-/over-scan, noise, dark frames, etc:


(1) Prescan problem - not flat. Overscan is flat.

(2) Dark current pattern: 50 e- per 2 hours ripple pattern and bad current at frame-transfer boundary - very hot; as much as 10,000 ADU in 2 hours. None of these problems were present in the PFIS detector. Some indication that the problem resides in the DSP code, though it is still under investigation.

(3) Software is stable and in place since VI phase for several months

(4) All modes are available except for drift scanning. The problem lies with coarse SDSU clock timing of 800 microsec res. Dave Carter is working on solution proposed by Ted. The video and slot modes need intensive in-situ testing (modes work in lab but not tested on the telescope).


Recent upgrades:

  • modifications for ACSI mode complete change in communication with controller from optical fibre to serial port.

  • addition of guider probe graphics superimposed on image.

  • fine-tuning of telescope positioning using moouse clicks on image.


On-line quality control:

  • images must be copied from SALTICAM disk for on-line analysis and transmission for intensive CPU stage.

  • tests with 1 Gb network successful. Max. size image copied in 3s.


Network:

  • 256 kb/s from Sutherland -> international, 256 kb/s from SAAO -> international

1 Mb/s from Sutherland to SAAO and to Telekom beachhead. First phase May 24, second by Jun 14. Infinite bandwidth from beach-head -> international, but when ?

  • the current 256 kb/s precludes video-conferencing.


Pipeline reductions:

CCD calibration software developed



  • aperture and weighted aperture photometry will use DoPhot, DAO-Phot and Allard’s method for CCD photometry.



12.PFIS Status Report (KN)


Integration:

Some parts are outstanding, including the polarizing beamsplitter, third (high resolution) etalon, remaining 24 interference filters.


Shipped:

30 Mar CCD detector and 05 Apr for the main instrument which was received on 14 Apr with no damage.



Fixes & retesting:

  • baffling

  • flexure retest

  • polarimetric testing

  • acceptance and characterization

  • polychromatic IQ tests with 12.5 micron pinholes yields 20-40 micron rms diameter (worst), 15-30 micron rms diameter (mean) where 1 arcsec = 117 microns, i.e. of order 0.2 arcsec between 420 and 780 nm.

  • CCD amplifier cross-talk is few x 10-3

  • Optical ghosts at 5 x 10-5 or less with an IF - worst case. These are specs of brightness of brightness part of the ghost relative to the integrated brightness of the PSF within 1.3 arcsec.

  • VPH efficiency - very close to specifications. Impressive. Note that manufacturer estimated delivered performance looked substantially off.

  • Focus runs - appears to be wave-dependent as expected, independent of gratings. HOWEVER, filters are high, and independent of wavelength because of thickness. Redoing all? No, but most will be made right. Luckily Pilot Group built in twice range

requested.0.3 arcsec slit, grating resolution is still slit-limited.

  • Reduction pipeline: based on the Gemini package and renamed "SALT":

CCD overscan, bias, crosstalk, mosaic (to 1/3 pixel, 0.04 arcsec); spectroscopy: cut out MOS spectra, wavelegnth cal, extract. Use spectroscopy (dispersion solution) to refine mosaic astrometry to milliarcsec level.

  • Central wave repeatability to 0.2 pixel.

  • Fabry-Perot: lower efficiency in the blue (60%); meets or above spec at 730 nm and red. Efficiency at 75-80% of FPRD.


Estimated CTC (cost-to-completion):

Currently ~$152k over cap (3.4%). Expect $225k over cap (4.5%) – linear extrapolation of CTC now to expected completion date.


Arrival in Sutherland

2 month slip of delivery to Cape Town since last meeting.

3 month " to Sutherland "
Schedule risk: detector firmware models; cold box flexure retest polarizing beamsplitter

Commissioning risks: availability of calibration system; training of PFIS SA; operational maturity; integration of commissioning and remaining telescope and payload tasks; hold PFIS in CT if necessary; characterize and train.
First week in June is "hold" to determine if PFIS goes to Sutherland. Will delay if telescope/payload is not ready.
VPH spectroscopy starts in mid-July; shared risk in late July or August.

Spares: $100k estimate


Additional costs:

4 electronic boxes - $25k

mechanism and mechanism simulators - $9.5k

- mechanisms are spares

- simulators to debug electronics

$34k capital

$58k salary
Spare control system: offer to do for fixed price at $102K

Another $5-7K in spare opto-mechanics.


Discussion on spares:

MB – spares needed now or can wait? KN: Cost-effective to do it now with people who have done it once.

RF – is it really needed? Who else does this? (DB, MB: NOT and NOAO)
Motion-1:

"The SSWG recommends that the Board makes provision for a PFIS initial spares budget, including a duplicate control system, and in a timeous manner." (DB/TW, carried unanimously).




13.PFIS NIR beam concept (Andy Sheinis)


CoDR review at next SSWG. Marcia Wolfe has been hired for this program as Instrument Scientist, while Mike Smith will be Project Engineer. Harlan Epps will do optical design.
Design philosophy:

  • Minimize risk

  • extensive performance modeling up front.

  • band-limit at the detector - semi-warm spectrograph and warm pupil with no internal pupil relay

  • cool camera (not cryo) - warm pupil

  • experienced VPH partner (UNC)

  • experienced F-P etalon partner (Rutgers)

  • existing instrumental frame

  • re-use mechanical design (articulation, mounts)


Constraints:

  • visible design, shared collimator

  • thermal constraints

  • weight budget

  • dispersion (collimator diameter and camera angle) all transmissive VPH – or at least transmissive gratings) with max R ~ 10,000

First-guess’ system parameters:



  • blue cut: 850 nm (optical design constraints)

  • red cut: 1.4-1.7 microns depends on thermal modeling

  • FOV: 8 arcmin

  • single 2kx2k detector

  • plate scale: 6 pix/arcsec = 108 micron/sec

  • camera ELF: 302 mm

  • beamsize: 149 mm

  • final f-ratio f/2.02

  • no cold stop, low emissivity warm stop

  • cooled camera (-20 to -80C)

  • band-limited detector

Compare to existing or planned systems it is unique in field, dual simultaneous optical-NIR and spectropolarimetry.


Performance Modeling: biggest risk. Note no limiting magnitude yet for design. Requires complete optical design to do this – in particular to mode scattered light.
Slit is worst "offender" in terms of contributions to thermal radiation. Solution is to coat slits with gold, but then there is a problem with scattered light. The solution here is a narcissus mirror, which blocks reflected light from coated slit coming in on either side of the collimator optics and self-reflects detector off potentially warm components in the beam.
What about moving baffle -- which is not anodized ? 2nd worst offender for background: warm enclosure radiation making it through in direct (0th order) path through the grating to the camera.

Again, solution is a narcissus mirror -- but this must articulate with grating. Telecentricity of camera is important.


Hope to get real stray light analysis complete for specific design by next meeting. Sample performance: will use Modtran-IV for atmosphere with full radiative transfer of upper atmosphere.
What resolution is sky set at? Consider narrow band pre-filter with low-res FP (tuneable filter)

and then disperse. This will significantly lower background.


Optical design:

A good design achieved w/o BaF2. Optical spot diagrams have 9-13m rms.


Grating design:

Does not want anamorphic factors .


Fabry-Perot:

Queensgate and Michigan as possible vendors. Wants to do testing at UW.


Camera assembly and integration at UW.

Cost: with contingency of 35% is $3.35M nominal. Expect a 15% match of $510k ($120k of which is from UNC). Net proposal would be for $2.8M from elsewhere.


Discussion:

LWR: Dewar cooling? A: Double cryotiger.

RG: What about NSF? A: Turned down by University for MRI. Will go again for ATI in Fall, and again for MRI in Jan 2006.

RF: Does thermal emissivity calculation include moving baffle and going off telescope? KHN: Can coat moving baffle with gold.

MB: question of whether another narcissus mirror would be needed. A: Don't know yet.

WK: What is time scale for building instrument? A: 3-year project from August 2005. LWR: That seems incredibly optimistic.


Final note: will send a questionnaire to SSWG for science feedback.

BC: CoDR has to have full-up calculations of limiting magnitudes. Can't gauge interest until this is know.

RF: In principle UW building NIR is good, and NIR always has been interesting, but, seems like we are coming up to limitation of HET for H and K. Is this really interesting to only do z and J?

MB: FP is very compelling because of limited band-pass.

DOD: In terms of SALTICAM experience, he presented several reports and then held a gun to the SSWG’s head and asked if it was wanted or not? We are interested in NIR, and Andy is interested in building it.

Price tag is high. Is it worth it? We must bear responsibility of giving him an answer soon because he needs to build something soon. Therefore process must be managed where there is good communication between AS and SSWG. We need more information and more backward and forward exchange of opinion. A set of questions need to be formulated, and then a set of expectations AS has of SSWG.

RF: Ill-defined interest since we don't know performance.

14.Report on HRS CDR (DB)


DB noted very competent external review panel present for the CDR.
Prior to the review, a trip was made to KiwiStar Optics (preferred optical fabricator), which operates under the umbrella of a former Government lab, which has a major share in current organization.


  • Comprehensive list of comments and suggestion to "tighten up" the design.

  • Recommendation regarding manufacture (optics).

  • Comments regarding mechanics.

  • Suggest to go for better radial velocity precision

The design is fundamentally sound, there are no perceived show-stoppers, and the project is in good shape technically.


On the basis of the review panel's report, the Project Scientist and Review Panel considers that the CDR has passed and that HRS is technically ready to go forward to construction.
Budget:

Following PDR: (Nov'04 baseline): $1.85M (Mar99) baseline budget including a 10% internal UC controlled (PI/PS) approved contingency. $185k project risk and $250k detector risk.

Following CDR: revised CTC increased $183k, i.e. revised cost of $2.122M (nominal) compared to $2.17 baseline (nominal).


  • devaluation of US$ against NZ$

  • increase capital expenditure

  • increased project management

  • new cost of Canterprise fee for contract negotiations (this last point is likely unacceptable for the Foundation)



15.HRS Status Report (Peter Cottrell)


Baseline design Jul 04

Revised design Nov 04

CDR Apr 05
Nod and shuffle for lowest resolution, unsliced mode only now, following recommendations and reconsideration.
Detectors: readout noise and dark current - concern that team may relax RN criteria. DC issue may be important if binning is significant (DOD estimates as much as 4 e- per hour per binned pixel).
Budget issue:

US$135k due to ROE

US$ 75k increase due to project and contract management.
Schedule:

CDR Apr 05

Integration Q2 06

ATP Q3 07


Contingency covers 100 days of schedule slip.
MB, DOD, KHN: note ATP slip relative to CDR date. Question of whether schedule is realistic.

Comments on specific CDR recommendations: cost & schedule implications need to be assessed and then propose whether to accept/defer.



  • active temperature control - will put in place

  • precision radial velocity mode - will consider ThAr injection and double-scrambling; future capability for iodine cell will be kept. Optimistic on throughput hit of double-fiber scrambler.

  • stacked stages at input

  • FPRD

  • detectors

  • test & verification



16.SALT Future Development & Gen 2 Instruments (DOD)


RF: advocacy of SALT-II

MB:


  • dilution is uncompetitive opening stance (dilution = death)

  • better to have less time on a more capable and competitive telescope.

  • Gen-2 instruments and SALT-II should not be mutually exclusive

  • mindful of developing successful match of VERY LIMITED pool of available / interested instrument builders in consortium and consortium science interests.

  • Cost threshold for SALT-II: $13-16M (depending on how cloned)

SALT30 (30-m) opportunity:



  • price of big science

  • window of opportunity

    • successful SALT

    • IAU symposium on big telescopes

RF: What are the science drivers that take us forward to define the planning?

KN: need mix of planning on different time-scales.

MB: technology development is important.


Will the SSWG support or recommend the following to the Board?

  • wide field capabilities: DOD size of field; cost

  • AO

  • NIR

There is a need for a process to move this development forward.


Both near term and long term goals and possibilities:

~2-3 years NIR/PFIS

~5 years A-O, SALT II

~12-15 years 20, 30, 50-m SALT


New funding: $2M from RAS (for HRS completion); >$1M from ANMH ?

17.First Science, First Light and Publicity Discussions


ACTION ITEM-2 (MAB): Produce list of images of galaxies for first light.

  • U and R band image.

  • try 380nm and U for galaxy to get [OII]3727 image.

ACTION ITEM-3 (DB): find out about press release distribution list of partner’s contacts.


Apparently there has been press release on the completion of the primary by SAAO, but there was no communication to partners.

Note: SAAO press-release on completing the telescope (Ken to check with Terry Devitt).

PAC: Special publication edition: MNRAS. Or common submission to astroph ?

18.SALT Partner Reports


UW: new VPH test lab

SAAO: meeting for local community in SA to develop SALT participation; solicitations for SALT TAC. DOD: to develop wide-field and AO concepts; plan for wide-field CCD camera on 74" in 18 months from now.

HET: Paper production at 2-3 per month. Texas hired two more extragal observers. Penn State hired brown-dwarf scientist. Adding tracker-focus software like SALT; should take 1.5-1.7 arcsec FWHM, which is now typical, to more like 1.1 arcsec FWHM. VIRUS under development.

Rutgers: Andrew Baker is a new faculty hire doing sub-mm (from Maryland), starting in year.

Pursuing drift-scan survey to support ACT. Putting in red dichroic: split >600 nm to PFIS blue-ward to SALTICAM.

CAMK: Contributed to SALTICAM software development (pipeline photometry processing).

CMU: Nothing to report.

UKSC: New positions (faculty and postdocs) within consortium; one chair.

Gott: E0.5M for buying SALTICAM; involved with building massively multiplexed IFU fed to 100's of spectrographs -- a large dark-energy project.

UC: nothing to report.




19.Summary of Action Items and Motions


Action items:
ACTION ITEM-1 (GS): update and distribute IQ error budget table.
ACTION ITEM-2 (MB): galaxy targets for First Light.

(DB): ask everyone for targets.


ACTION ITEM-3 (DB): find out about press release distribution.
Summary of Motions:
Motion-1:

"SSWG recommends to the Board make the provision of PFIS spare budget, including a duplicate control system, in a timeous manner."






Dostları ilə paylaş:
Orklarla döyüş:

Google Play'də əldə edin


Verilənlər bazası müəlliflik hüququ ilə müdafiə olunur ©muhaz.org 2017
rəhbərliyinə müraciət

    Ana səhifə